OpenStudy (anonymous):

Triangle ABC has side lengths 3, 4, and 5. Do the side lengths form a Pythagorean triple? Explain.

jimthompson5910 (jim_thompson5910):

If you can show that a^2 + b^2 = c^2 is true, where a,b,c are the sides of the triangle (c is the longest side), then a,b,c is a pythagorean triple and they form a right triangle.

jimthompson5910 (jim_thompson5910):

So in this case, a = 3, b = 4, c = 5

jimthompson5910 (jim_thompson5910):

Plug them into the equation and see if it holds true.

OpenStudy (anonymous):

@math_genius123

OpenStudy (math_genius123):

\[a ^{2}+b ^{2}=c ^{2}\]

OpenStudy (anonymous):

and are all non zero whole numbers ?

OpenStudy (math_genius123):

what do ya mean

OpenStudy (anonymous):

3,4, and 5 ?

OpenStudy (math_genius123):

yeah

OpenStudy (shamim):

3^2+4^2=5^2 is it correct

OpenStudy (shamim):

so it follow pythagorean triangle